# Rotational dynamics of a disk in a thin film of weakly nematic fluid subject to linear friction

### v1

@@ Abstract
Dynamics at low Reynolds numbers experiences recent revival in the fields of biophysics and active matter. While in bulk isotropic fluids it is exhaustively studied, this is less so in anisotropic fluids and in confined situations. Here, we combine the latter two by studying the rotation of a disk-like inclusion in a uniaxially anisotropy, globally oriented, incompressible two-dimensional fluid film. In terms of a perturbative expansion in parameters that quantify anisotropies in viscosity and in additional linear friction with a supporting substrate or other type of confinement, we derive analytical expressions for the resulting hydrodynamic flow and pressure fields as well as for the resistance and mobility coefficients of the rotating disk. It turns out that, in contrast to translational motion, the solutions remain well-behaved also in the absence of the additional linear friction. Comparison with results from finite-element simulations show very good agreement with those from our analytical calculations. Besides applications to describe technological systems, for instance, in the area of microfluidics and thin cells of aligned nematic liquid crystals, our solutions are important for quantitative theoretical approaches to fluid membranes and thin films in general featuring a preferred direction.
@@ I Introduction
Here, $\eta_{\mathrm{S}}$ is the two-dimensional membrane viscosity, and $\operatorname{Bq}=\eta_{\mathrm{S}}/\eta a$ is a dimensionless quantity known in interfacial rheology as the Boussinesq number [22], with $\eta$ denoting the three-dimensional viscosity of a bulk fluid surrounding the thin fluid layer. In addition, $\gamma$ denotes for Euler’s constant. The diffusion coefficients are then derived from the resistance coefficients using the classical Stokes-Einstein relations [23, 24].
@@
It is worth noting that the Saffman–Delbrück theory is applicable in the limit of large $\operatorname{Bq}$. The general case of a finite Boussinesq number was later examined [25, 26]. Based on hydrodynamic membrane models [20, 27], the viscosity of the lipid membrane can be determined by measuring both the rotational and translational diffusion coefficients of membrane-linked particles, along with their radii [28].
@@
Confining boundaries and dampening interfaces are known to reduce the drag coefficient of microparticles, leading to decreased diffusion [29, 30, 31]. Along these lines, effects of the system size on the rotational diffusion of membrane proteins have been explored using molecular dynamics simulations [32, 33, 34]. Observations include a decreased rotational diffusion coefficient when compared to an infinitely extend bulk system, influenced by hydrodynamic interactions with the confining boundaries [32]. Specifically, the rotational dynamics of a sample protein under cylindrical nanopore confinement has been examined [35]. The simulation study revealed a twofold reduction in the rotational diffusion coefficient compared to the bulk value when the confinement radius is twice the hydrodynamic radius of the protein. A versatile numerical scheme for predicting diffusion coefficients of arbitrarily shaped objects embedded in lipid bilayer membranes has also been presented [36].
@@
The rotational dynamics of membrane proteins have been investigated using molecular-dynamics simulations of coarse-grained membranes and two-dimensional Lennard-Jones fluids under different crowding conditions [37]. It has been observed that the Saffman–Delbrück theory [20] accurately describes the size-dependence of rotational diffusion under dilute conditions, whereas more complex scaling laws emerge under crowded conditions. More recently, the anisotropic diffusion of membrane proteins has been computationally investigated [38]. Generally, anisotropy can arise from the anisotropic shape of proteins and protein complexes or from the anisotropic interactions of proteins with the lipid bilayer [39, 40, 41].
@@
Beyond the focus of our previous work on translational motion [44], we now address rotational motion of the circular disk in a globally uniaxially ordered, incompressible two-dimensional fluid at low Reynolds numbers. As before, we include the influence of additional linear friction, which may be induced by a supporting substrate or other types of environment of the two-dimensional fluid film or membrane. To this end, we utilize a technique of two-dimensional Fourier transformation to solve the hydrodynamic equations for the velocity and pressure fields [47, 48, 49, 50, 51, 52]. In cases of weak anisotropy of both viscosity and of linear friction, we apply a perturbative approach up to second order in these small parameters. We derive approximate expressions for the hydrodynamic fields, resistance and mobility coefficients. Specifically and in contrast to the translational case [44], in the absence of additional linear friction we obtain exact analytical expressions in the form of convergent series expansions. Our theoretical predictions are supplemented by a comparison with results from finite-element simulations, showing very good agreement between the two different approaches.
@@ II Flow in two-dimensional anisotropic fluids with friction
wherein $\boldsymbol{\sigma}(\bm{r})=p(\bm{r})\,\bm{I}+\widetilde{\boldsymbol{\sigma}}(\bm{r})$ denotes the stress tensor. Here, $p(\bm{r})$ represents the pressure field, $\bm{I}$ stands for the unit tensor, and $\widetilde{\boldsymbol{\sigma}}(\bm{r})$ incorporates the effect of viscous dissipation into the dynamics. Moreover, $\bm{f}(\bm{r})$ denotes the two-dimensional inplane force density acting on the fluid and $\bm{M}$ represents the tensor of linear friction associated with the bottom and/or top environment. $\bm{v}(\bm{r})$ is the velocity field. We remark that the convention for the sign of $\widetilde{\boldsymbol{\sigma}}(\bm{r})$ can vary across different sources in the literature and, accordingly, the associated tensor might be defined with a minus sign in some cases.
@@
represent two dimensionless numbers quantifying the degree of anisotropy. Here, $\nu_{i}$, $i\in 1,\dots,5$, are five coefficients of viscosity for the case of uniaxial symmetry [57] of SI unit kg/s in two spatial dimensions. Additionally, for convenience, we introduce the abbreviations $\alpha_{\parallel}^{2}=m_{\parallel}^{2}/\nu_{3}$ and $\alpha_{\perp}^{2}=m_{\perp}^{2}/\nu_{3}$, respectively. We note that both $\alpha_{\parallel}^{2}$ and $\alpha_{\perp}^{2}$ possess dimensions of inverse length squared.
@@ III Hydrodynamic flow fields
Next, we set $\bm{f}_{2n+1}=\left(ic_{2n+1},-d_{2n+1}\right)^{\mathrm{T}}$ with $(\cdot)^{\top}$ denoting the transpose. These coefficients are chosen as real numbers. Their values are determined based on the boundary conditions prescribed at the surface of the disk, as detailed below. To ensure the coefficients $c_{2n+1}$ and $d_{2n+1}$ are real-valued, we additionally require that $c_{-(2n+1)}=-c_{2n+1}$ and $d_{-(2n+1)}=d_{2n+1}$. Therefore, the force density given by Eq. (14) can be expressed as
@@
with $I_{2n+1}$ and $K_{2n+1}$ denoting the $(2n+1)$th-order modified Bessel functions of the first and second kind, respectively. Here, $\mathcal{P}\Delta_{mn}$ denotes the principal part of the series expansion of $\Delta_{mn}$ around $\sigma=0$. Specifically, the principal part corresponds to the part of the Laurent series expansion of $\Delta_{mn}$ characterized by negative exponents [61]. This result is valid only when $w\leq 1$, which corresponds to the physical scenario of the fluid area outside the disk. A detailed analytical solution of a similar integral is provided in Appendix B of Ref. 44.
@@ 10 $\checkmark$
At leading orders in $A$ and $\beta$, we find that only the second and fourth Fourier modes in the radial and azimuthal velocities, along with the zeroth mode of azimuthal velocity, remain non-vanishing. Additionally, the sixth Fourier mode for pressure persists. Notably, the zeroth-order mode of the pressure vanishes. When accounting for second-order terms, the sixth and eighth Fourier modes for the velocity as well as the tenth Fourier mode for the pressure remain non-vanishing. We summarize the non-vanishing Fourier modes of the velocity in the multi-Taylor expansion in parameters $A$ and $\beta$ in Tab. 3, while we present the corresponding results for the pressure in Tab. 3.
@@
$\bm{v}$ $=a\Omega\left(\bm{v}^{\mathrm{IB}}+\bm{v}^{*}\right)\,,$ (35a)
$p$ $=\frac{\Omega\nu_{3}}{a}\left(p^{\mathrm{IB}}+p^{*}\right)\,,$ (35b)
wherein $\bm{v}^{\mathrm{IB}}$ is the solution for an isotropic Brinkman fluid for $A=\beta=b=0$. Specifically, $\bm{v}^{*}$ denotes the leading-order correction due to uniaxial anisotropy in viscosity and friction. Defining the abbreviation
$\bm{q}=\big{(}K_{0}(c),K_{1}(c),K_{2}(c),K_{0}(r),K_{1}(r),K_{2}(r)\big{)}^{\top}\,,$ (36)
the solution for the velocity and pressure field for an isotropic Brinkman fluid is obtained as
$v_{\rho}^{\mathrm{IB}}=0\,,\qquad v_{\theta}^{\mathrm{IB}}=\frac{q_{5}}{q_{2}}\,,\qquad p^{\mathrm{IB}}=0\,.$ (37)
We remark that $q_{2}=(c/2)\left(q_{3}-q_{1}\right)$ and $q_{5}=(r/2)\left(q_{6}-q_{4}\right)$.
@@
$p^{*}=\frac{bc}{2}\frac{q_{6}}{q_{2}}\,\sin(2\theta)\,.$ (38)
Unlike the case of translational motion considered previously [44], for which the solution diverges logarithmically as $c\to 0$, it remains well-behaved in this limit for rotational motion. The non-vanishing Fourier modes for the velocity and pressure fields are obtained up to second order in $A$ as presented in Tab. 4. We infer that the coefficients $u_{n}$ defining the Fourier series for the pressure field vanish for odd $n$ in the case of $b=0$.
@@ IV Resistance coefficient
$R^{\mathrm{IB}}=2\pi c\,\xi_{2}$ (47)
to denote the resistance coefficient in an isotropic two-dimensional Brinkman fluid medium. This result is in full agreement with Ref. 63, which reports the translational and rotational drag coefficients for a disk moving in a liquid membrane supported by a rigid substrate. Additionally, if $\alpha_{\parallel}=\alpha_{\perp}=0$, we recover the classical result for the rotational resistance coefficient as given by Eq. (1). The leading-order corrections to the resistance coefficient due to the anisotropies in viscosity $A$ and in friction $\beta$ are given by
@@
$\mu$ $=\frac{\mu^{\mathrm{IB}}}{\nu_{3}}\Big{(}1+\mu^{(1,0)}A+\mu^{(0,1)}\beta$
@@
where $\mu^{\mathrm{IB}}=1/R^{\mathrm{IB}}$ represents the rotational mobility within an isotropic two-dimensional Brinkman fluid medium. The correction terms arising from the anisotropies in viscosity and friction are derived to leading order in $A$ and $\beta$ as
$\mu^{(1,0)}$ $=-\frac{R^{(1,0)}}{R^{\mathrm{IB}}}\,,$ (52a)
$\mu^{(0,1)}$ $=-\frac{R^{(0,1)}}{R^{\mathrm{IB}}}\,.$ (52b)
@@
$\mu^{(2,0)}$ $=\left(\mu^{(1,0)}\right)^{2}-\frac{R^{(2,0)}}{R^{\mathrm{IB}}}\,,$ (53a)
$\mu^{(1,1)}$ $=2\mu^{(1,0)}\mu^{(0,1)}-\frac{R^{(1,1)}}{R^{\mathrm{IB}}}\,,$ (53b)
$\mu^{(0,2)}$ $=\left(\mu^{(0,1)}\right)^{2}-\frac{R^{(0,2)}}{R^{\mathrm{IB}}}\,.$ (53c)
@@
In Fig. 2 we present a comparison of the correction to the scaled rotational mobility function for different combinations of viscosity and friction anisotropy, derived from our perturbative analytical theory (lines) and fully resolved finite-element simulations (symbols), see Sec. VI. Results are plotted against the scaled friction coefficient $c=\alpha_{\parallel}c$. Generally, the corrections are very small compared to the bulk mobility $\mu^{\mathrm{IB}}$, making them challenging to determine in real situations. For $(A,\beta)=(0.5,0)$, representing the case of viscosity anisotropy only, the correction reaches its maximum value as $c\to 0$, exhibiting a logistic-like evolution, and vanishes as $c\to\infty$, corresponding to the limit of infinite friction. For $(A,\beta)=(0,0.5)$, representing the case of friction anisotropy only, the curve exhibits a bell-shaped form, peaking around $c\sim 1$. In this case, the correction is positive across the entire range of $c$, indicating that friction anisotropy enhances the rotational motion of the disk. By combining both modes of anisotropy, for $(A,\beta)=(0.5,0.5)$, an interesting evolution is observed, showcasing the complex interplay between viscosity anisotropy and friction anisotropy. Overall, there is excellent agreement between the theoretical predictions and the simulation results.

### v2

@@ Abstract
Dynamics at low Reynolds numbers experiences recent revival in the fields of biophysics and active matter. While in bulk isotropic fluids it is exhaustively studied, this is less so in anisotropic fluids and in confined situations. Here, we combine the latter two by studying the rotation of a disk-like inclusion in a uniaxially anisotropic, globally oriented, incompressible two-dimensional fluid film. In terms of a perturbative expansion in parameters that quantify anisotropies in viscosity and in additional linear friction with a supporting substrate or other type of confinement, we derive analytical expressions for the resulting hydrodynamic flow and pressure fields as well as for the resistance and mobility coefficients of the rotating disk. It turns out that, in contrast to translational motion, the solutions remain well-behaved also in the absence of the additional linear friction. Comparison with results from finite-element simulations show very good agreement with those from our analytical calculations. Besides applications to describe technological systems, for instance, in the area of microfluidics and thin cells of aligned nematic liquid crystals, our solutions are important for quantitative theoretical approaches to fluid membranes and thin films in general featuring a preferred direction.
@@ I Introduction
Here, $\eta_{\mathrm{S}}$ is the two-dimensional membrane viscosity, and $\operatorname{Bq}=\eta_{\mathrm{S}}/\left(\eta a\right)$ is a dimensionless quantity known in interfacial rheology as the Boussinesq number [22], with $\eta$ denoting the three-dimensional viscosity of a bulk fluid surrounding the thin fluid layer. In addition, $\gamma$ denotes Euler’s constant. The diffusion coefficients are then derived from the resistance coefficients using the classical Stokes–Einstein relations [23, 24].
@@
It is worth noting that the Saffman–Delbrück theory is applicable in the limit of large $\operatorname{Bq}$. The general case of a finite Boussinesq number was later examined [25, 26]. Based on hydrodynamic membrane models [20, 27], the viscosity of the lipid membrane can be determined by measuring both the rotational and translational diffusion coefficients of membrane-linked particles, along with their radii [28]. In these analyses, the membrane is treated as a thin liquid layer of finite thickness.
A simple phenomenological approach has later been introduced by Evans and Sackmann [29] to characterize the weak dynamic coupling of a fluid membrane to an adjacent solid substrate. Utilizing this framework, the inertialess equations of motion for membrane flow have been solved for the steady translation and rotation of a disk-like particle. Remarkably, the resulting drag coefficients show functional dependencies on the dimensionless particle size that are very similar to those derived for particle motion in a membrane bounded by semi-infinite fluid domains. Yet, the scaling with the particle size differs [25].
According to the Evans–Sackmann model, the translational and rotational resistance coefficients for a disk-like inclusion of radius $a$ in the membrane were determined in a scaled form as
$\frac{R_{\mathrm{T}}}{4\pi\eta_{\mathrm{S}}}=\frac{c^{2}}{4}+c\,\frac{K_{1}(c)}{K_{0}(c)}\,,\quad\frac{R_{\mathrm{R}}}{4\pi a^{2}\eta_{\mathrm{S}}}=1+\frac{c}{2}\frac{K_{0}(c)}{K_{1}(c)}\,,$ (2)
where $c$ is a dimensionless parameter defined as $c=a\left(b_{\mathrm{S}}/\eta_{\mathrm{S}}\right)^{1/2}$ and $b_{\mathrm{S}}$ marks an intrinsic coefficient of friction between the membrane and the substrate. Expanding the right-hand sides of Eqs. (2) in a power series of $c$ up to order $c^{2}$ yields Eq. (1), which is derived using the Saffman–Delbrück model, for $c=2/\operatorname{Bq}$. Our contribution relates to the Evans–Sackmann model due to its simplicity and practicality, when compared to the original Saffman–Delbrück model, particularly in the presence of a substrate.
@@
Confining boundaries and dampening interfaces are known to reduce the mobility coefficient of microparticles, leading to decreased diffusion [30, 31, 32]. Along these lines, effects of the system size on the rotational diffusion of membrane proteins have been explored using molecular dynamics simulations [33, 34, 35]. Observations include a decreased rotational diffusion coefficient when compared to an infinitely extended bulk system, influenced by hydrodynamic interactions with the confining boundaries [33]. Specifically, the rotational dynamics of a sample protein under cylindrical nanopore confinement has been examined [36]. The simulation study revealed a twofold reduction in the rotational diffusion coefficient compared to the bulk value when the confinement radius is twice the hydrodynamic radius of the protein. A versatile numerical scheme for predicting diffusion coefficients of arbitrarily shaped objects embedded in lipid bilayer membranes has also been presented [37].
@@
The rotational dynamics of membrane proteins have been investigated using molecular-dynamics simulations of coarse-grained membranes and two-dimensional Lennard-Jones fluids under different crowding conditions [38]. It has been observed that the Saffman–Delbrück theory [20] accurately describes the size-dependence of rotational diffusion under dilute conditions, whereas more complex scaling laws emerge under crowded conditions. The study reports a crowding-induced transition between Saffman–Delbrück-like and stronger scanlings, where the rotational diffusion coefficient $D_{\mathrm{R}}\sim a^{-\nu}$ with $\nu>2$. In systems of increasing levels of crowding, the Saffman–Delbrück model fits the data for small disks, while larger disks exhibit a stronger scaling of $\nu=3$. More recently, the anisotropic diffusion of membrane proteins has been computationally investigated [39]. Generally, anisotropy can arise from the anisotropic shape of proteins and protein complexes or from the anisotropic interactions of proteins with the lipid bilayer [40, 41, 42].
@@
Beyond the focus of our previous work on translational motion [45], we now address rotational motion of the circular disk in a globally uniaxially ordered, incompressible two-dimensional fluid at low Reynolds numbers. As before, we include the influence of additional linear friction, which may be induced by a supporting substrate or other types of environment of the two-dimensional fluid film or membrane. We note in this context that in the previous translational case the additional linear friction has a stabilizing effect for the system and the associated mathematical solution. In an infinitely extended two-dimensional fluid, the mathematical solution of the Stokes equation for the flow resulting from the net translational motion of an object immersed under no-slip conditions diverges logarithmically as a function of the distance from that object. This is referred to as the Stokes paradox. Additional linear friction stabilizes the solution and cancels the divergence. Yet, such an effect is absent in the case of rotational motion, so that additional linear friction does not take the central role of stabilization. Generally for the flow field, the leading order of dependence on the distance from the object for its rotational motion is obtained by derivation of the leading order for its translational motion [48, 49, 50, 51]. Therefore, we find decay with inverse distance instead of logarithmic divergence in the rotational case.
More in detail, we here utilize a technique of two-dimensional Fourier transformation to solve the hydrodynamic equations for the velocity and pressure fields [52, 53, 54, 55, 56, 57]. In cases of weak anisotropy of both viscosity and of linear friction, we apply a perturbative approach up to second order in these small parameters. We derive approximate expressions for the hydrodynamic fields, resistance and mobility coefficients. Specifically and in contrast to the translational case [45], in the absence of additional linear friction we obtain exact analytical expressions in the form of convergent series expansions. Our theoretical predictions are supplemented by a comparison with results from finite-element simulations, showing very good agreement between the two different approaches.
@@ II Flow in two-dimensional anisotropic fluids with friction
wherein $\boldsymbol{\sigma}(\bm{r})=p(\bm{r})\,\bm{I}+\widetilde{\boldsymbol{\sigma}}(\bm{r})$ denotes the stress tensor. Here, $p(\bm{r})$ represents the pressure field, $\bm{I}$ stands for the unit tensor, and $\widetilde{\boldsymbol{\sigma}}(\bm{r})$ incorporates the effect of viscous dissipation into the dynamics. Moreover, $\bm{f}(\bm{r})$ denotes the two-dimensional inplane force density acting on the fluid and $\bm{M}$ represents the tensor of linear friction associated with the bottom and/or top environment. $\bm{v}(\bm{r})$ is the velocity field.
@@
In this situation, the stress tensor $\bm{\widetilde{\sigma}}(\bm{r})$ is reduced to account only for the pure dissipative stress arising from gradients in the fluid flow [62, 63],
$\widetilde{\sigma}_{ij}=\sigma_{ij}^{\text{D}}={}-\nu_{ijkl}\nabla_{l}v_{k}\,.$ (4)
In this expression, $\nu_{ijkl}$ refer to the components of the viscosity tensor of uniaxial symmetry [62],
$\nu_{ijkl}$ $=\nu_{2}(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})+\bar{\nu}n_{i}n_{j}n_{k}n_{l}+(\nu_{4}-\nu_{2})\delta_{ij}\delta_{kl}$
$\quad+(\nu_{3}-\nu_{2})(n_{i}n_{k}\delta_{jl}+n_{i}n_{l}\delta_{jk}+n_{j}n_{k}\delta_{il}+n_{j}n_{l}\delta_{ik})$
$\quad+(\nu_{5}-\nu_{4}+\nu_{2})(\delta_{ij}n_{k}n_{l}+\delta_{kl}n_{i}n_{j})\,,$ (5)
where $\nu_{i}$ ($i\in 1,\dots,5$) are five coefficients of viscosity for the case of uniaxial symmetry [62]. Here, we introduced the abbreviation
$\bar{\nu}=2\left(\nu_{1}+\nu_{2}-2\nu_{3}\right)\,.$ (6)
In the case of an isotropic fluid, the first term on the right-hand side of Eq. (5) of coefficient $\nu_{2}$ refers to the shear viscosity and thus here is related to the viscosity $\eta_{\mathrm{S}}$. The third term of coefficient $\nu_{4}-\nu_{2}$ relates to volume/bulk (or area in two dimensions) viscosity. All other terms are absent for an isotropic fluid as they explicitly involve the role of the director $\bm{\hat{n}}$. We remark that the convention for the sign of $\widetilde{\boldsymbol{\sigma}}(\bm{r})$ can vary across different sources in the literature and, accordingly, the associated tensor might be defined with a minus sign in some cases.
@@
represent two dimensionless numbers related to anisotropy. Additionally, for convenience, we introduce the abbreviations $\alpha_{\parallel}^{2}=m_{\parallel}^{2}/\nu_{3}$ and $\alpha_{\perp}^{2}=m_{\perp}^{2}/\nu_{3}$, respectively. We note that both $\alpha_{\parallel}^{2}$ and $\alpha_{\perp}^{2}$ possess dimensions of inverse length squared.
@@ III Hydrodynamic flow fields
Next, we set $\bm{f}_{2n+1}=\left(ic_{2n+1},-d_{2n+1}\right)^{{\top}}$ with $(\cdot)^{\top}$ denoting the transpose. These coefficients are chosen as real numbers. Their values are determined based on the boundary conditions prescribed at the surface of the disk, as detailed below. To ensure the coefficients $c_{2n+1}$ and $d_{2n+1}$ are real-valued, we additionally require that $c_{-(2n+1)}=-c_{2n+1}$ and $d_{-(2n+1)}=d_{2n+1}$. Therefore, the force density given by Eq. (18) can be expressed as
@@
with $I_{2n+1}$ and $K_{2n+1}$ denoting the $(2n+1)$th-order modified Bessel functions of the first and second kind, respectively. Here, $\mathcal{P}\Delta_{mn}$ denotes the principal part of the series expansion of $\Delta_{mn}$ around $\sigma=0$. Specifically, the principal part corresponds to the part of the Laurent series expansion of $\Delta_{mn}$ characterized by negative exponents [66]. This result is valid only when $w\leq 1$, which corresponds to the physical scenario of the fluid area outside the disk. A detailed analytical solution of a similar integral is provided in Appendix B of Ref. 45.
@@ 10 $\checkmark$
At leading orders in $A$ and $\beta$, we find that only the second and fourth Fourier modes in the radial and azimuthal velocities, along with the zeroth mode of azimuthal velocity, remain non-vanishing. Additionally, the sixth Fourier mode for pressure persists. Notably, the zeroth-order mode of the pressure vanishes. When accounting for second-order terms, the sixth and eighth Fourier modes for the velocity as well as the tenth Fourier mode for the pressure remain non-vanishing. We summarize the non-vanishing Fourier modes of the velocity expressed in terms of $g_{n}$ and $h_{n}$ as given in the form of a multi-Taylor expansion in parameters $A$ and $\beta$ in Tab. 3, while we present the corresponding results for the pressure expressed in terms of $u_{n}$ in Tab. 3.
The non-vanishing Fourier modes for the velocity and pressure fields are obtained up to second order in $A$ as presented in Tab. 4. We infer that the coefficients $u_{n}$ defining the Fourier series for the pressure field vanish for odd $n$ in the case of $b=0$.
@@
$\bm{v}$ $=a\Omega\left(\bm{v}^{\mathrm{IES}}+\bm{v}^{*}\right)\,,$ (39a)
$p$ $=\frac{\Omega\nu_{3}}{a}\left(p^{\mathrm{IES}}+p^{*}\right)\,,$ (39b)
wherein $\bm{v}^{\mathrm{IES}}$ is the solution for the isotropic Evans–Sackmann model implying $A=\beta=b=0$. Specifically, $\bm{v}^{*}$ denotes the leading-order correction due to uniaxial anisotropy in viscosity and friction. The solution for the velocity and pressure field for the isotropic Evans–Sackmann model is obtained as
$v_{\rho}^{\mathrm{IES}}=0\,,\qquad v_{\theta}^{\mathrm{IES}}={\frac{K_{1}(r)}{K_{1}(c)}}\,,\qquad p^{\mathrm{IES}}=0\,.$ (40)
@@
$p^{*}=\frac{bc}{2}{\frac{K_{2}(r)}{K_{1}(c)}}\,\sin(2\theta)\,.$ (41)
Unlike the case of translational motion considered previously [45], for which the solution diverges logarithmically as $c\to 0$, it remains well-behaved in this limit for rotational motion. Accordingly, the solution for rotational motion remains well-behaved in the absence of friction. Moreover, from Eq. (10), it follows that for $A=\beta=b=0$ the dimensionless viscosity coefficients become $\lambda_{\pm}=1$, leading to vanishing pressure. This is also predicted by the Evans–Sackmann theory, for which $v_{\rho}=0$ and $v_{\theta}$ is a function solely of $\rho$.
@@ IV Resistance coefficient
$R^{\mathrm{IES}}=2\pi c\,\xi_{2}$ (50)
to denote the resistance coefficient for the isotropic Evans–Sackmann model. This result is in full agreement with Ref. 29, which reports the translational and rotational drag coefficients for a disk moving in a fluid membrane supported by a rigid substrate. Additionally, if $\alpha_{\parallel}=\alpha_{\perp}=0$, we recover the classical result for the rotational resistance coefficient as given by Eq. (1). The leading-order corrections to the resistance coefficient due to the anisotropies in viscosity $A$ and in friction $\beta$ are given by
@@
$\mu$ $=\frac{\mu^{\mathrm{IES}}}{\nu_{3}}\Big{(}1+\mu^{(1,0)}A+\mu^{(0,1)}\beta$
@@
where $\mu^{\mathrm{IES}}=1/R^{\mathrm{IES}}$ represents the rotational mobility for the isotropic Evans–Sackmann model. The correction terms arising from the anisotropies in viscosity and friction are derived to leading order in $A$ and $\beta$ as
$\mu^{(1,0)}$ $=-\frac{R^{(1,0)}}{R^{\mathrm{IES}}}\,,$ (55a)
$\mu^{(0,1)}$ $=-\frac{R^{(0,1)}}{R^{\mathrm{IES}}}\,.$ (55b)
@@
$\mu^{(2,0)}$ $=\left(\mu^{(1,0)}\right)^{2}-\frac{R^{(2,0)}}{R^{\mathrm{IES}}}\,,$ (56a)
$\mu^{(1,1)}$ $=2\mu^{(1,0)}\mu^{(0,1)}-\frac{R^{(1,1)}}{R^{\mathrm{IES}}}\,,$ (56b)
$\mu^{(0,2)}$ $=\left(\mu^{(0,1)}\right)^{2}-\frac{R^{(0,2)}}{R^{\mathrm{IES}}}\,.$ (56c)
@@
In Fig. 2 we present a comparison of the correction to the scaled rotational mobility function for different combinations of viscosity and friction anisotropy, derived from our perturbative analytical theory (lines) and fully resolved finite-element simulations (symbols), see Sec. VI. Results are plotted against the scaled friction coefficient $c=\alpha_{\parallel}a$. Generally, the corrections are very small compared to the bulk isotropic mobility $\mu^{\mathrm{IES}}$, making them challenging to determine in real situations. For $(A,\beta)=(0.5,0)$, representing the case of viscosity anisotropy only, the correction reaches its maximum value as $c\to 0$, exhibiting a logistic-like evolution, and vanishes as $c\to\infty$, corresponding to the limit of infinite friction. For $(A,\beta)=(0,0.5)$, representing the case of friction anisotropy only, the curve exhibits a bell-shaped form, peaking around $c\sim 1$. In this case, the correction is positive across the entire range of $c$, indicating that friction anisotropy enhances the rotational motion of the disk. By combining both modes of anisotropy, for $(A,\beta)=(0.5,0.5)$, an interesting evolution is observed, showcasing the complex interplay between viscosity anisotropy and friction anisotropy. Overall, there is excellent agreement between the theoretical predictions and the simulation results.
We remark that in an infinitely extended, isotropic two-dimensional fluid medium, there is no coupling between translational and rotational components in the resistance matrix of a moving disk. Similarly, we do not find such coupling for a disk translating in a uniaxially anisotropic fluid. That is, the translation of the disk does not induce rotational motion and vice versa due to this lack of coupling.